OpenStudy (christos):

f(x) = (2x +1)^3 f'(x) = 6(2x + 1)^2 f''(x) = 48x + 24 I need to know when its concave up/down increasing /decreasing and the inflection points I am new to this kind of stuff

OpenStudy (rulnick):

First derivative is nonnegative for all real x, so f is non-decreasing. Second derivative is everywhere matching the sign of x+1/2, so there is an inflection point at x=-1/2. The function is concave down on x<-1/2 and concave up on x>-1/2.

OpenStudy (christos):

how did you find the -1/2

OpenStudy (rulnick):

f''(x)=0 at x=-1/2

OpenStudy (christos):

Ok and something more are my derivative calculations correct? f(x) = (2x +1)^3 f'(x) = 6(2x + 1)^2

OpenStudy (rulnick):

Yes, all were perfect!

OpenStudy (christos):

so its not decreasing that means its always increasing? Kinda what's the interval?

OpenStudy (christos):

(0,infinity) increasing?

OpenStudy (rulnick):

non-decreasing means increasing or flat. it is flat at the inflection point, increasing everywhere else

OpenStudy (rulnick):

so increasing on the entire real line except at -1/2, where it is flat (deriv=0)

OpenStudy (christos):

here it asks me the open interval on which f is increasing what should I put? (-inf,-1/2)U(-1/2,int) ?

OpenStudy (rulnick):

yes, very nicely done

OpenStudy (christos):

and decreasing interval*

OpenStudy (rulnick):

empty set

OpenStudy (christos):

like I just say "it's not decreasing anywhere" ?

OpenStudy (rulnick):

yes

OpenStudy (christos):

Alright, thank you!

OpenStudy (rulnick):

welcome